OpenStudy (anonymous):

What is the value of the coefficient "b" when the quadratic equation y = (3x − 5)(2x + 3) is written in standard form? 19 6 −15 −1

OpenStudy (anonymous):

do you know how to multiply this out?

OpenStudy (anonymous):

no can you help me

OpenStudy (anonymous):

you have to do 4 multiplications \[ (3x − 5)(2x + 3)\] \[=3x\times 2x+3\times 3x-5\times 2x-5\times 3\] is a start

OpenStudy (anonymous):

sometimes called "foil" by mediocre math teachers, but it is really the distributive law we have more work to do

OpenStudy (anonymous):

\[=3x\times 2x+3\times 3x-5\times 2x-5\times 3=6x^2+9x-10x-15\] then combine like terms and get \[6x^2-x-15\]

OpenStudy (anonymous):

now if we call \[6x^2-x-15\] as \[ax^2+bx+c\] we see that \(a=6,b=-1,c=-15\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

all steps were there, hope they are clear i didn't skip any so you can look this over if you like

OpenStudy (anonymous):

so the answer is -15